# YITP-24-22

### v1

@@ Abstract
Discussion – In this Letter, we proposed a novel probe of primordial NG, specifically targeting the non-linearity parameter $\tau_{\rm NL}$ associated with the four-point correlation of the primordial curvature perturbation. This probe is powerful for the previously unconstrained ultra-light-mass window for PBHs, as the NG on small scales is sensitively captured by the spatial clustering of ultra-light PBHs, which leads to an enhanced curvature perturbation on larger scales after those PBHs dominate the universe and generate GWs at the time of PBH evaporation whose amplitude can be large enough to be detected. Accordingly, a distinctive bi-peak structure can emerge in the GW spectrum, with the amplitude of the secondary GW peak at low frequencies being directly related to $\tau_{\rm NL}$. Notably, by accounting for upper bounds on the GW amplitude derived from existing theoretical and observational constraints, we establish for the first time to the best of our knowledge a joint constraint on the correlations between the effective non-linearity parameter $\bar{\tau}_{\rm NL}$ and the mass and abundance of ultra-light PBHs, which can effectively yield an upper bound on primordial NGs.
The proposed probe has several immediate cosmological implications. At the observational level, there is a promising chance for the GW signal induced by our mechanism to fall into the frequency bands of the forthcoming GW experiments including LISA, ET, SKA and BBO, giving thus the potential detection of the aforementioned distinctive spectral shape in near future. At the theoretical level, one can test via the above GW portal any feasible inflation models that give rise to ultra-light PBHs with scale-dependent NGs. Additionally, it is worth mentioning that the novel NG probe introduced here can in principle be extended to any phenomenological scenario involving compact objects with a monochromatic or extended mass function Papanikolaou (2022), coupled with a transient eMD epoch. Those aforementioned subjects deserve to be studied in detail, and we leave related explorations to future work.

### v2

@@ Abstract
In Fig. 2 we show the upper bound constraint on $\bar{\tau}_{\mathrm{NL}}\mathcal{P}_{\cal R}(k)$ (color bar axis) as a function of the PBH mass $M_{\mathrm{PBH}}$ ($x$-axis) and the initial PBH abundance at formation $\Omega_{\mathrm{PBH,f}}$ ($y$-axis). The BBN bound on $\Omega_{\mathrm{PBH,f}}$, Eq. (13), excludes the upper right dark gray region. The quadrilateral pattern on the left side of the color map and the corresponding gap around $M_{\text{PBH}}\sim 10^{6}\rm{g}$ arise from the constraints imposed by the LIGO-VIRGO-KAGRA (LVK) experiment while the vertical grey region in the mass range $[10^{8}\mathrm{g},10^{9}\mathrm{g}]$ is excluded from the effect of Hawking evaporation of PBHs in this mass range on the abundances of light elements produced during BBN Boccia et al. (2024). We also require that PBHs dominate before they evaporate. This leads to the lower bound on the $\Omega_{\mathrm{PBH,f}}$, $\Omega_{\mathrm{PBH,f}}>6\times 10^{-10}\frac{10^{4}\mathrm{g}}{M_{\mathrm{PBH}}}$ Domènech et al. (2021a), which excludes the lower left light grey region.
Discussion – In this Letter, we propose a novel probe of primordial NG, specifically targeting the non-linearity parameter $\tau_{\rm NL}$. This probe is powerful for the previously unconstrained ultra-light-mass window for PBHs, as primordial NG on small scales is sensitively captured by the spatial clustering of PBHs. Remarkably, we find a distinctive bi-peaked GW spectral shape, with the amplitude of the secondary GW peak at low frequencies being directly related to $\tau_{\rm NL}$. Notably, by accounting for upper bounds on the GW amplitude derived from existing theoretical and observational constraints, we establish for the first time to the best of our knowledge a joint constraint on the correlations between the effective non-linearity parameter $\bar{\tau}_{\rm NL}$ and the mass and abundance of ultra-light PBHs, which can effectively yield an upper bound on primordial NGs.
The proposed probe has several immediate cosmological implications. At the observational level, there is a promising chance for the GW signal induced by our mechanism to fall into the frequency detection bands of forthcoming GW experiments, pointing towards the potential detectability of the aforementioned distinctive GW signal in the near future. At the theoretical level, one can test via the above GW portal any feasible inflation models that give rise to ultra-light PBHs with scale-dependent NGs. Additionally, it is worth mentioning that the novel NG probe introduced here can in principle be extended to any phenomenological scenario involving compact objects with a monochromatic or extended mass function Papanikolaou (2022), coupled with a transient eMD epoch. Those aforementioned subjects deserve to be studied in detail, and we leave related explorations to future work.
Finally, we need to highlight that future measurements and/or upper bound constraints on $\Delta N_{\mathrm{eff}}$ at the percent level in the coming years will set tighter constraints on local-type primordial non-Gaussianities on small scales through our PBH induced GW portal whereas forecasts on local-type non-Gaussianities from DESI Krolewski et al. (2024); Rezaie et al. (2024) and CMB-S4 Sohn and Fergusson (2019) on larger scales (namely the LSS and CMB ones) will act as a complementary to our GW portal non-Gaussianity probe.